OpenStudy (anonymous):

How would I find the zeros of the equation x^3 - 5x^2 + 6x = 0 ?

terenzreignz (terenzreignz):

Factor.

OpenStudy (anonymous):

How? Like x(x^2 - 5x + 6) ?

terenzreignz (terenzreignz):

Yes... that's good. Actually, the other part (x^2 - 5x + 6) could still be factored further ;)

OpenStudy (anonymous):

Really? I don't know, I was never good at factoring. How would that work, would you mind trying to explain it?

terenzreignz (terenzreignz):

Crud... I'm not good at explaining how to factor... what two numbers, when multiplied, give 6 and when added, give -5?

OpenStudy (anonymous):

-3 and -2, right?

terenzreignz (terenzreignz):

yup. So.. the factors are...?

terenzreignz (terenzreignz):

x-3 and x-2 try multiplying them ^_^

OpenStudy (anonymous):

Ah... this is why I hate factoring... So it would be like this, right? x(x^2 - 5x + 6) = 0 x(x-2)(x-3) ?

terenzreignz (terenzreignz):

That's right. Can you proceed from there? x(x-2)(x-3) = 0

OpenStudy (anonymous):

x = 0, 2, 3 or would it be x = 0, -2, -3

terenzreignz (terenzreignz):

Why don't you find out? Plug them in, see which ones work...

OpenStudy (anonymous):

Okay, I'll try~

terenzreignz (terenzreignz):

Okay, I'll help you out. When you have a set of factors that are equal to zero, it suffices for just one of the factors to be zero. So since you have x(x-2)(x-3) = 0 then just solve x = 0 x-2 = 0 x-3 = 0 And you'll get your answers ^_^

OpenStudy (anonymous):

So it would be 0, 2, and 3 then right?

OpenStudy (anonymous):

Well, thanks for your help! It was very helpful so I think I understand better~

terenzreignz (terenzreignz):

No problem ^_^

OpenStudy (anonymous):

Would that still work the same if the problem had the first with a power of 4 and not 3?

terenzreignz (terenzreignz):

Well, the factoring would be different, wouldn't it?

terenzreignz (terenzreignz):

Unless all of them have powers increased by 1...